### v1

# Evolution of boundedly rational learning in games

Abstract
People make strategic decisions multiple times a day. We act strategically in negotiations, when we coordinate our actions with others, or when we choose with whom to cooperate. The resulting dynamics can be studied with evolutionary game theory. This framework explores how people adapt their decisions over time, in light of how effective their strategies have proven to be. A crucial quantity in respective models is the strength of selection. This quantity regulates how likely individuals switch to a better strategy when given the choice. The larger the selection strength, the more biased is the learning process in favor of strategies with large payoffs. Therefore, this quantity is often interpreted as a measure of rationality. Traditionally, most models take selection strength to be a fixed parameter. Instead, here we allow the individuals’ strategies and their selection strength to co-evolve. The endpoints of this co-evolutionary process depend on the strategic interaction in place. In many prisoner’s dilemmas, selection strength increases indefinitely, as one may expect. However, in snowdrift or stag-hunt games, it can either converge to a finite value, or we observe evolutionary branching altogether – such that different individuals arrive at different selection strengths. Overall, this work sheds light on how evolution might shape learning mechanisms for social behavior. It suggests that boundedly rational learning is not only a by-product of cognitive constraints. Instead it might also evolve as a means to gain strategic advantages.
Keywords: evolutionary game theory |||| adaptive dynamics |||| learning |||| social dilemma
Introduction
People constantly make strategic decisions. These decisions range from rather inconsequential choices (e.g., whether to do a small favor) to major commitments (e.g., choosing a business partner). The process by which individuals arrive at their decisions can be analyzed with evolutionary game theory [1, 2, 3]. This framework treats individuals as players who interact in games. They can choose among different strategies and win a payoff which depends both on their own decisions and those of their co-players. Importantly, evolutionary game theory does not require individuals to choose optimal decisions from the outset. Rather, individuals are assumed to engage in the same interaction (or sufficiently similar ones) many times. Over the course of these interactions, they adapt their behavior to yield higher payoffs. The respective dynamical process allows for several interpretations. We can think of strategies as being genetically determined and evolving across generations [4, 5]. Alternatively, models of learning consider how individuals switch between strategies within their lifetime [6, 7, 8, 9, 10, 11, 12, 13]. Typically, these two interpretations are studied separately or interchangeably. In practice, however, the two approaches are intertwined: social behavior is often shaped by learning, while learning mechanisms are shaped by evolutionary forces [14, 15, 16, 17].
Herein, we ask how evolutionary forces might shape a key parameter of many stochastic learning models, the selection strength or intensity of selection. This parameter $\beta\!\geq\!0$ determines to which extent an individual’s choice to switch strategies is driven by payoff differences  (Fig. 1A). When this parameter is small, players are largely insensitive to their strategies’ payoffs. They mostly switch from one strategy to another by chance, with payoff differences only slightly affecting that probability (Fig. 1B, left). The respective limit of weak selection ($\beta\!\rightarrow\!0$) has become a major paradigm in evolutionary game theory, partly because it simplifies many calculations [18, 19, 20, 21, 22]. In this limit, learning takes the form of an (almost unbiased) random walk on the space of all strategies. In contrast, as selection strength becomes large, players only switch strategies if the new strategy is superior (Fig. 1B, right). The respective limit of strong selection ($\beta\!\rightarrow\!\infty$) has become another major paradigm to study evolutionary games [23, 24, 25]. As one advantage of that limit, respective results are closely aligned with key concepts in classical game theory. For example, in this limit strict Nash equilibria correspond to absorbing states of the learning process [26, 27]. Herein, we thus interpret the value of $\beta$ as a measure of how rational individuals are when updating their strategies [similar to the interpretation of a related quantity in the ‘logit equilibrium’ 28, 29, 30]. Accordingly, we think of finite $\beta$ values as indicating boundedly rational learning.
Figure 1: An overview of the model. A, In stochastic evolutionary game theory, individuals continually get a chance to revise their strategy. We model this revision process with introspection dynamics [31, 32, 33, 34, 35, 36, 37]. Here, a player compares its current strategy to a randomly chosen alternative. If $\pi$ is the payoff of the current strategy, and $\pi^{\prime}$ the payoff of the alternative, the player’s switching probability is given by a Fermi function $\phi(\pi^{\prime}\!-\!\pi)$. This function captures that players are more likely to switch if the alternative strategy is more beneficial. B, The exact shape of the switching probability depends on a parameter $\beta$, the strength of selection. For low $\beta$, payoff differences between strategies become irrelevant. Here, strategy changes are mostly driven by chance. As $\beta$ becomes larger, individuals increasingly favor those alternative strategies with high payoffs. Thus, $\beta$ is often interpreted as a measure of the players’ rationality or their learning effectiveness. We are interested in the evolution of this parameter. Our process unfolds on two timescales. C, In the short run, each player’s selection strength $\beta$ is fixed. Given their selection strength, they choose between different strategies in a stage game. The depicted example shows a round in which player 1 switches from strategy C (cooperation) to strategy D (defection). We iterate this learning process for many rounds. Based on these iterations, we compute the average frequency of each possible game outcome (illustrated as a black-and-white gradient). This allows us to compute the players’ expected payoff as a function of their $\beta$ values. D, In the long-run, we let the players’ selection strength $\beta$ evolve. We model this long-run process with adaptive dynamics [38, 39, 40, 41].
Except for a handful of studies discussed further below [42, 43], most models interpret selection strength as a fixed quantity, which is the same for everyone. Here, we relax this assumption. We address two related questions. First, we are interested in learning processes among individuals with fixed but different selection strengths. We refer to individuals with larger $\beta$ as more effective learners, and those with lower $\beta$ as less effective learners. We ask to which extent more effective learners are able to systematically outperform their opponents, by choosing their strategies more optimally.
Second, we explore the dynamics when an individual’s selection strength is itself an evolving trait. To this end, we use a framework akin to the indirect evolutionary approach used to study the evolution of preference [44, 45, 46, 47, 48]. The dynamics unfolds on two time scales. In the short run, the players’ selection strengths are fixed. Players interact in a given game with others, adopt new strategies based on their selection strength, and as a result, earn a certain average payoff (Fig. 1C). In the long run, we allow the selection strength to evolve, based on the average payoffs the respective players yield. To the extent that selection strength is an inheritable trait, this assumption may reflect that more effective learners have more offspring (Fig. 1D). We explore this long-run dynamics analytically and with simulations, based on the adaptive dynamics formalism [38, 39, 40, 41]. To get an intuitive understanding, we focus in the main text on the most simple class of games: symmetric games among two players with two strategies. These games include, for example, the prisoner’s dilemma, the stag-hunt game, or the snowdrift game. However, we also show that the respective intuitions carry over to more general multiplayer games.
Once a player’s selection strength is an evolvable trait, one might expect an evolutionary dynamics towards ever-increasing selection strengths. After all, players with larger selection strength are better at identifying strategies with high payoffs. Hence, such players should have a fitness advantage. Remarkably, however we do not find such a runaway dynamics for all games. Instead in many snowdrift games, selection strength converges to a finite value. Moreover, in many stag-hunt games we observe evolutionary branching. Here, an initially monomorphic population splits into two or more subpopulations, each with a different selection strength. Interestingly, we obtain these results even without assuming that more effective learners need to exert more cognitive effort and hence pay some complexity cost. Instead, we obtain these results because an ineffective learning process can turn out to be a strategic advantage. In this way, our study contributes to a growing literature exploring how strategic considerations can explain behaviors that are seemingly at odds with individually rational decision-making [44, 45, 46, 47, 48].
Model and results
Our model assumes a separation of timescales. In the short run, players have fixed but possibly different selection strengths. They repeatedly engage in a given one-shot game, and they learn to adopt more profitable strategies over time. We model this short-run learning process with introspection dynamics, a process particularly useful to study games with some asymmetry [31, 32, 33, 34, 35, 36, 37]. In the long run, the players’ selection strengths are allowed to evolve. We describe this long-run process with adaptive dynamics [38, 39, 40, 41], a classical tool to study the evolution of continuous traits. In the following, we describe the two processes in turn.
A description of the short-run dynamics
To explore learning among players with different selection intensities, we consider simple $2\times 2$ games. That is, we consider two players who can choose among two possible strategies, $\mathbf{C}$ and $\mathbf{D}$ (we think of the two strategies as cooperation and defection, but this interpretation is irrelevant for our results). The choices of the two players determine their payoffs. Payoffs are assumed to be symmetric, and given by the (normalized) payoff matrix
$\begin{array}[]{c|cc}&\mathbf{\;C\;}&\mathbf{\;D\;}\\ \hline\cr\mathbf{\;C\;}&1&-B\\ \mathbf{\;D\;}&1-A&0\\ \end{array}$ (1)
We refer to this matrix as the stage game. As usual, player 1 chooses a row of this matrix, player 2 chooses a column, and the entries represent the payoffs of player 1. Depending on the signs of $A$ and $B$, we recover four major game classes: prisoner’s dilemma ($A\!<\!0$, $B\!>\!0$), stag-hunt game ($A\!>\!0$, $B\!>\!0$), snowdrift game ($A\!<\!0$, $B\!<\!0$), and harmony game ($A\!>\!0$, $B\!<\!0)$.
To describe how players learn to adopt more profitable strategies over time, we use introspection dynamics [31, 32]. This process assumes that the two individuals interact in the same stage game over many time steps. Each time step, a randomly chosen player is given a chance to revise its strategy. That player then compares its current payoff $\pi$ with the payoff $\pi^{\prime}$ the player could have obtained by choosing the other strategy (keeping the co-player’s strategy fixed). As illustrated in Fig. 1A, the player switches to the other strategy with a probability that depends on the payoff difference [6, 7, 8],
$\phi=\frac{1}{1+{\rm e}^{-\beta_{i}\,(\pi^{\prime}-\pi)}}.$ (2)
The parameter $\beta_{i}\!\geq\!0$ represents player $i$’s selection strength. For small $\beta_{i}$, this function approaches $1/2$ irrespective of the payoff difference (Fig. 1B left). In this limit, players decide whether to switch strategies based on a coin toss. For large $\beta_{i}$, the function approaches a step function (Fig. 1B right). Here, the player always switches if the alternative strategy yields the better payoff. We iterate this basic updating process for many time steps. As a result, players repeatedly get a chance to switch their strategies  (Fig. 1C, bottom). Based on this process, we record how frequently we observe each of the four possible game outcomes over time. In the right panel of Fig. 1C, we illustrate these overall frequencies using a black-and-white gradient. A darker shading indicates game outcome that are observed more often.
Introspection dynamics is particularly convenient because the expected frequencies of the four game outcomes can be computed analytically. To this end, we represent the learning dynamics as a Markov chain (see Supporting Information for details). The Markov chain has four states, corresponding to the four possible outcomes in each time step, CC, CD, DC, DD (the first letter refers to the first player’s strategy, and the second letter to the second’s). For finite $\beta_{1},\beta_{2}$, this Markov chain has a unique stationary distribution $\mathbf{u}(A,B,\beta_{1},\beta_{2})\leavevmode\nobreak\ :=\leavevmode\nobreak\ (u_{\textbf{CC}},u_{\textbf{CD}},u_{\textbf{DC}},u_{\textbf{DD}})$. This distribution captures how often we observe each of the four game outcomes on average, over the course of many rounds. Once we have the stationary distribution, it becomes straightforward to compute player 1’s average payoff, by multiplying the four frequencies $u_{ij}$ with the respective stage game payoffs,
$\Pi_{\beta_{2}}(\beta_{1})=1\cdot u_{\textbf{CC}}-B\cdot u_{\textbf{CD}}+(1-\!A)\cdot u_{\textbf{DC}}+0\cdot u_{\textbf{DD}}.$ (3)
The respective results become particularly simple when both players have the same selection strength, $\beta_{1}\!=\!\beta_{2}$. In that case, one can show that the stationary distribution simplifies to
$\mathbf{u}(A,B,\beta,\beta)=\frac{1}{2+e^{A\beta}+e^{B\beta}}(e^{A\beta},1,1,e^{B\beta}).$ (4)
For weak selection ($\beta\!\rightarrow\!0$), this formula implies that all game outcomes are observed equally often, $\mathbf{u}=(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4})$. For strong selection ($\beta\!\rightarrow\!\infty$), the stationary distribution depends on the game parameters $A$ and $B$. For example, for the prisoner’s dilemma, (4) implies that the two players defect in almost all rounds ($u_{\textbf{DD}}\!=\!1$); in the snowdrift game, they anti-coordinate ($u_{\textbf{CD}}\!=\!u_{\textbf{DC}}\!=\!\frac{1}{2}$); and in the stag-hunt game they choose the strategy that is risk-dominant [$u_{\textbf{DD}}\!=\!1$ if $A\!<\!B$, $u_{\textbf{CC}}\!=\!1$ if $A>B$, see Ref. 31].
Using the same formalism, we can also explore the case when players have different selection strengths. In the Supporting Information, we derive a general expression of the stationary distribution $\mathbf{u}(A,B,\beta_{1},\beta_{2})$ as a function of $\beta_{1}$ and $\beta_{2}$, which is now more complex. In the following, we use that expression to explore whether a more effective learner (one with larger $\beta$) would gain a payoff advantage in various stage games.
In Fig. 2A–D, we consider a prisoner’s dilemma. We assume the two players can either have a small or a comparably large selection strength, $\beta\!\in\!\{2,5\}$. For a symmetric situation in which both players use the same $\beta$, we find that players learn to defect more often when $\beta$ is large (Fig. 2A,B). This result is intuitive: the larger the selection strength, the better players are at identifying the individually optimal strategy of defecting. Once players are asymmetric, it is the more effective learner who defects more often (Fig. 2C). As a result, this player ends up with a better average payoff.
Figure 2: Introspection dynamics among players with different selection strengths. To illustrate the impact of selection strength on the learning dynamics, we consider two stage games, a prisoner’s dilemma (A–C) and a snowdrift game (D–F). In each case, we assume selection strengths may assume two possible values. They can either be comparably low (bright avatar) or high (dark avatar). For all possible combinations of selection strengths, we depict how often the respective players would obtain one of the four possible payoffs of the respective stage game. Based on this stationary distribution, we compute the players’ expected payoffs. For the prisoner’s dilemma we observe that the higher selection strength dominates the lower selection strength (panel D). In contrast, for the snowdrift game, it is the low-selection strength player who gains the larger average payoff (panel H).
In Fig. 2E–H, we consider a snowdrift game. For illustration, the possible selection strengths are now $\beta\!\in\!\{5,15\}$. In the symmetric case among players with equal selection strength, a large $\beta$ again makes it more likely that players implement a Nash equilibrium (here, CD or DC). In the depicted example, they coordinate on a Nash equilibrium in $74\%$ of all rounds when $\beta\!=\!5$, but in $90\%$ of all rounds when $\beta\!=\!15$. Remarkably, however, once players have different selection strengths, it is the less effective learner who obtains the larger average payoff. By acting myopically against a comparably rational opponent, such players ensure that their more preferred equilibrium is visited more often (CD in Fig. 2G). In this way, they obtain a superior payoff – $\Pi_{15}(5)\!=\!0.86>0.52\!=\!\Pi_{5}(15)$ – despite being less sophisticated in their strategy choices.
Based on these results, we can construct a ‘super-game’ matrix (Fig. 2D,H). This matrix represents the interaction of two learners who choose between different selection strengths. The entries of the matrix depict the learners’ average payoffs over the course of the learning process (as derived in the previous panels). For the prisoner’s dilemma, the super-game indicates that each learner always prefers to have a higher selection strength. In the snowdrift game example, we observe the opposite. Here, it is the lower selection strength that dominates the higher one.
Overall, these two examples provide a proof of principle. They illustrate the non-trivial effects of asymmetric selection strengths. These observations call for a more systematic analysis, considering more games and a larger set of possible selection strengths. We provide such an analysis in Fig. S1 of the Supporting Information. In addition, the above results raise the question whether selection strength would necessarily move towards ever-increasing values if it was an evolving trait. We explore that issue in the next section.
A description of the long-run dynamics
To explore how the players’ selection strengths $\beta$ might change over time, we use adaptive dynamics [38, 39, 40]. That is, we consider a large and well-mixed population of players (Fig. 1D). Each player randomly meets other population members to interact in a series of pairwise games. As a result, they obtain a payoff as described in the previous section on introspection dynamics. We define a player’s fitness to be its average payoff across all its interaction partners. Players with larger fitness are more likely to reproduce. The offspring of a reproducing player inherits the parent’s trait $\beta$ with probability $1\!-\!\mu$. With the complementary probability $\mu$, the offspring’s selection strength is slightly mutated, by adding some small noise to the parent trait. Overall, these assumptions define a dynamical process on the population level, with $\beta$ being the evolving trait. This process is straightforward to simulate. We give a detailed description of the algorithm in the Material and Methods, and provide the respective code online.
Figure 3: Adaptive dynamics of selection strength for three different social dilemmas. We explore the adaptive dynamics for three different social dilemmas: the prisoner’s dilemma, the snowdrift game, and the stag-hunt game. For each case, we depict a realization of an individual-based simulation (upper panels) and a pairwise invasibility plot (lower panels). Each simulation is initialized at a monomorphic population with a selection strength of $\beta\!=\!0$. Over time, mutations introduce additional selection strength values. Individuals obtain payoffs by randomly interacting with other population members, and learning strategies through introspection dynamics. They reproduce according to their payoff. For details, see Material and Methods. Pairwise invasibility plots illustrate the dynamics assuming a monomorphic resident population. They display (in color) which rare mutant traits ($y$-axis) have a positive invasion fitness $s_{x}(y)$, given the current resident ($x$-axis). Colored regions above the diagonal indicate that mutants with a higher trait value than the resident are favored to invade. Colored regions below the diagonal indicate evolution towards smaller trait values. Dashed lines indicate the position of singular trait values. A,B, For the depicted prisoner’s dilemma, we observe a dynamics of ever-increasing selection strengths. C,D, In the snowdrift game, the evolving $\beta$ values converge to a finite value. E,F, In the stag-hunt game, we observe evolutionary branching. After this occurs, there are two subpopulations. Members of the first subpopulation have low $\beta$ values, and hence they choose strategies mostly at random. Members of the second subpopulation exhibit high $\beta$ values; they tend to best respond to their respective opponent.
Initially, we assume that the population is monomorphic, with a selection strength of $\beta\!=\!0$. To gain some intuition of the further dynamics, we revisit the previous two examples. Fig. 3A shows simulations for the prisoner’s dilemma. Here, we observe that $\beta$ evolves towards ever-increasing values. This outcome is consistent with the naive expectation that more effective learners should have a fitness advantage, and consistent with our analysis in Fig. 2A–D. For the other example, the snowdrift game, simulations are shown in Fig. 3C. Here, even though larger selection strengths are initially favored, the dynamics stops at some value well below $\beta\!=\!6$. That is, in contrast to our naive expectation – but consistent with our analysis in Fig. 2E–H – we observe the evolution of boundedly rational learning.
We can gain some analytical understanding of these results by considering the respective pairwise invasibility plots (PIPs, Fig. 3B,D). These plots depict the direction of evolution as the population size becomes infinitely large, and as mutations become sufficiently rare. In that case, we can assume the population to be monomorphic, such that all players have the same (resident) trait of $\beta\!=\!x$. To explore whether a mutant with a different trait $y\!\neq\!x$ might invade, we compute the mutant’s invasion fitness,
$s_{x}(y):=\Pi_{x}(y)-\Pi_{x}(x).$ (5)
In this formula, $\Pi_{x}(y)$ is the payoff of a mutant with selection strength $\beta\!=\!y$ in a resident population with $\beta\!=\!x$. This payoff is given by (3). The mutant can selectively invade the resident if and only if this invasion fitness is positive. In pairwise invasibility plots, the respective region for which $s_{x}(y)\!>\!0$ is highlighted in color. For the prisoner’s dilemma, this colored area comprises the entire space above the main diagonal (Fig. 3B). This means that independent of the exact selection strength $x$ of the resident, a mutant with trait $y$ can invade if and only if $y\!>\!x$. This implies the evolution of ever-increasing selection strengths.
In contrast, in the snowdrift game in Fig. 3D, mutants with a higher selection strength are only favored when the resident’s selection strength is sufficiently small (if $x\!<\!x^{*}$ for some threshold $x^{*}$). We can numerically compute this threshold by setting the so-called selection gradient equal to zero (see Material and Methods). For the depicted example, we obtain $x^{*}\!=\!5.58$, in line with the evolutionary simulations. The pairwise invasibility plot also indicates that this singular point is evolutionarily stable. Once the population settles at a selection strength of $\beta\!=\!x^{*}$, neither a more effective nor a less effective learner can invade (as indicated by the white neighborhood around the vertical dashed line).
Interestingly, these two examples of an ever-increasing selection strength and an evolutionarily stable finite selection strength are not the only possibilities. Instead, Fig. 3E shows a third scenario, using a stag-hunt game. Here, we observe evolutionary branching. At first, the monomorphic population again approaches a singular point, in this case at $x^{*}\!=\!3.01$. Once there, both mutant types can invade, those that are more effective learners and those that are less effective learners (indicated by the colored neighborhood of the vertical dashed line in Fig. 3F). This is the fingerprint of a branching point. From here, the population starts diverging in opposite directions. There is a sub-population for which the selection strength increases, and another one for which it decreases. The subsequent dynamics is more complex to describe analytically; we do so in the Supporting Information and Fig. S2. In the main text, we provide some intuition for why branching occurs in the first place. To this end, it is again instructive to look at the stationary distribution of introspection dynamics, for two different selection strengths $\beta$ (see Fig. S3A–D). When both players use the same $\beta$, a larger $\beta$ makes players choose the less efficient DD equilibrium more often (because defection is risk-dominant). If one player instead has a low $\beta$, both players occasionally coordinate on the more efficient CC equilibrium, while often avoiding harmful mis-coordination on the CD-outcome.
Once selection strength is allowed to evolve, its dynamics also has important implications on the population’s average payoff. For example, in the prisoner’s dilemma, as individuals become more effective in their learning process, they increasingly learn to adopt the individually optimal strategy of defection. As a result, overall payoffs decrease in time (Fig. S4, left column). For the other two examples, the long-run effects on payoffs are more advantageous. Both for the snowdrift and the stag-hunt game, we observe an initially negative but eventually positive trend in the players’ payoffs (Fig. S4, middle and right column). In particular, the evolutionary branching in the stag-hunt game turns out to be both individually and collectively beneficial.
So far we have studied the evolution of selection strength for a few isolated $2\!\times\!2$ games. To identify more general patterns, in a next step we systematically vary the parameters $A$ and $B$ of the stage game. For each resulting game, we compute the numerically exact position of the singular point (if it exists), and its stability properties. The result is Fig. 4. Here, white regions correspond to cases without a singular point. For the corresponding stage games, we predict selection strengths to increase indefinitely. Colored regions correspond to cases with a finite singular point. Here, we distinguish orange regions (where the point is evolutionarily stable) and purple regions (where we observe branching). Based on this figure, we make three observations.
Figure 4: Evolutionary dynamics across all social dilemmas. So far, we described three special instances of games; here we systematically repeat this analysis for all $2\!\times\!2$ games. For each pair of game parameters $A$ and $B$, we numerically check whether there exists a finite singular trait value $\beta^{*}$. If it exists, we explore whether the respective trait value is evolutionarily stable (orange) or a branching point (purple). The color gradient represents the position of the singular trait: the lighter, the larger the value of $\beta^{*}$. Regions without a finite singular trait are left white. Black symbols indicate the position of the three examples displayed in Fig. 3.
First, even though most prisoner’s dilemmas lead to ever-increasing selection strengths (as in Fig. 3A), there are exceptions. These exceptions occur if the absolute value $|A|$ is small (if there is only a small temptation to defect when the co-player cooperates). Here, we observe evolution towards a finite value of $\beta$. The underlying mechanism is different from the mechanism in the snowdrift game discussed earlier. Because $|A|$ is small compared to $B$, a co-player is more likely to occasionally switch to cooperation if the focal player cooperates. This positive influence on the co-player’s cooperation rate can make it worthwhile to have a finite $\beta$, and to cooperate occasionally (Fig. S3E–H).
Second, in the snowdrift game, we observe evolution towards a finite selection strength for all games with $B\!>\!A$. In addition, as we consider a wider range of parameter values than considered in Fig. 4, we find another region exhibiting evolution towards finite $\beta$, see Fig. S5. This second region is given by $B\!<\!A\!-\!1$; we provide a derivation of these conditions in the Supporting Information. In both regions, the mechanism favoring finite selection strengths is the same: By having lower selection strengths, learners ensure they spend more time in their individually preferred equilibrium.
Finally, in the stag-hunt game, we only observe evolutionary branching when it is risk-dominant to defect (that is, for $A\!<\!B$). If instead cooperation is risk-dominant, the prevalence of CC increases as a function of the players’ selection strength. Mutual cooperation is the optimal outcome for both players; hence, an ever-increasing selection strength is individually and mutually beneficial.
Beyond pairwise games
Above, we documented several cases with an unexpected dynamics (e.g., evolution towards finite selection strengths or evolutionary branching). In each case, the dynamics occurred because at least one player gained some strategic advantage from being a less effective learner. In general, however, one might expect that these strategic advantages are strongest in pairwise games. Among two players, one player’s (possibly irrational) behavior tends to have a comparably strong effect on how the co-player might react. This effect could wane in groups of size $N\!>\!2$. To further explore the robustness of our results, in the Supporting Information we study the dynamics of three classes of public goods games. The three classes differ in the shape of the public good’s benefit function. Benefits are either linear, they show diminishing returns [49], or they take the form of a threshold function [50, 51]. While the computational steps involved in analyzing the adaptive dynamics are the same as before, the mathematical expressions now become more intricate. Importantly, however, the qualitative dynamics are similar to before. Depending on the payoff function, again we recover the three previously observed cases of indefinite evolution, convergence to a finite evolutionarily stable state, or evolutionary branching (Fig. S6, for $N\!=\!3$).
Discussion
Evolutionary game theory explores how strategies and traits change over time [1, 2, 3]. Such an evolutionary approach can be motivated in two ways. On the one hand, the respective dynamical equations may describe long-run biological evolution. Here, individuals with better strategies are assumed to reproduce more often [4, 5]. On the other hand, evolutionary models can also be interpreted as a representation of an individual’s short-run learning process. Such models assume that better strategies are more readily adopted [6, 7, 8, 9, 10, 11, 12]. Herein, we blend these two interpretations. We ask how a long-term biological process may shape the effectivity of individual learning. To this end, we study the evolution of a key quantity in stochastic learning models, the players’ selection strength. This quantity determines to which extent payoff advantages sway individuals to change their strategies. When selection strength is small, individuals tend to ignore their strategies’ payoffs [18, 19, 20, 21, 22]. Here, strategy choices are mostly determined by chance. As the selection strength becomes larger, individuals become increasingly sensitive to (even minor) payoff differences.
Motivated by these observations, herein we interpret selection strength as an individual trait that measures how rational individuals act when updating their strategies. We use this perspective to study the evolution of boundedly rational learning in games. To this end, we consider a model with two time scales. In the short run, individuals have fixed but possibly different selection strengths. We explore the resulting learning processes with introspection dynamics [31, 32, 33, 34, 35, 36, 37]; for the $2\times 2$ games we focus on, this process is equivalent to the logit-response dynamics studied in economics [32, 52]. We use this short-term perspective to ask whether individuals with larger selection strengths are able to take advantage of their opponents.
In the long run, we also allow the individuals’ selection strength to evolve. Here, naively one might expect ever-increasing selection strengths. After all, large selection strengths make an individual’s learning process more effective, allowing them to better distinguish profitable strategies from unprofitable ones. Surprisingly, however, we identify several important game classes in which selection strength converges towards a finite value. In some cases, ineffective learning seems to act as a commitment device. By learning more erratically, individuals commit themselves to use strategies that run counter to their immediate interests. In such a case, more effectively learning opponents may be swayed to react, and to coordinate on an equilibrium that eventually benefits the more ineffective learner. These findings have interesting parallels in the biology literature on the red-king effect [53, 54, 55]. That literature suggests that in mutualisms among two co-evolving species too, it is the slower evolving species that eventually has an evolutionary advantage.
More generally, our study contributes to a research area that explores the origins and consequences of boundedly rational behavior. Gigerenzer and Selten point out that there are two complementary directions to the study of bounded rationality, “psychological plausibility” and “ecological rationality” [56]. The first direction focuses on analyzing the mind (such as people’s cognition or their emotions) to make sense of behavior. The second direction suggests to look at the structure of environments to understand particular heuristics and cognitive biases that are domain-specific to those environments [57]. Our contribution relates to this second direction. We vary an individual’s social ecology (the specific stage game under consideration) and let a trait associated with rational learning evolve. Based on this setup, we explore how distinct environments shape individual adaptations.
Our model is also directly related to two interesting studies by Szabó, Vukov, and Szolnoki [42, 43]. They also consider a model in which two traits co-evolve. These traits are the individuals’ strategies, and a noise parameter that is inversely related to our selection strength. Our work deviates from these studies in many ways. For example, we look at well-mixed populations while they study behavior on a regular lattice; we treat selection strength as a continuous variable while they consider a discrete set of possible noise parameters; we explore the entire range of $2\!\times\!2$ games whereas they consider a one-dimensional subspace of games. Most importantly, however, they consider simulations in which strategies and the noise parameter evolve at the same rate. Instead, we assume a separation of timescales. As a result, they do not address most of the questions that are central to our approach. For example, because they do not allow for lasting differences in an individual’s learning capabilities, they do not explore to which extent more effective learners are able to outperform their opponents. This question plays a crucial role in our setup, and it drives all of our subsequent results (such as the emergence of evolutionary branching, which they do not observe).
Naturally, there are several ways how one could make our model more complex. For example, one could go beyond the symmetric one-shot games studied herein. Such work can shed light on the effects of asymmetric payoffs [33, 34, 35], repeated games [58], or heterogenous environments [59, 60]. In particular, such an approach could help us understand how people have developed psychological mechanisms to cope with the variability of situations they face. Another interesting direction would be to study the evolution of boundedly rational learning in structured rather than well-mixed populations. A framework of adaptive dynamics in such populations is readily available [61, 62]. This literature shows that network structure strongly affects the way individuals cooperate [63, 64, 65]. With a framework similar to ours, one could explore how network structure affects the way individuals learn. This work could investigate, for example, whether heterogeneous networks promote heterogeneity in people’s learning mechanisms. Finally, our results can also find future applications in multi-agent reinforcement learning. Specifically, one could conceive extensions where agents dynamically adapt their learning rates as a function of the dilemmas faced [66].
Materials and Methods
We provide a detailed account of both the short-run learning dynamics and the long-run evolutionary dynamics in the Supporting Information. In the following we briefly sketch our methods for the long-run evolutionary dynamics. For that, we use a combination of individual-based simulations and a theoretical analysis in the infinite population limit.
Evolutionary simulations
Our evolutionary simulations use a similar setup as Döbeli et al [67]. We start with a monomorphic population with trait $\beta=0$. At each time step, two individuals, $i$ and $j$, are randomly chosen from the population. These two individuals then compete for reproduction. To model this competition, each of the two individuals independently engage in a sample of three encounters with other random individuals. For each encounter, their payoff is computed with (3). By averaging over the three realizations, we compute each player’s average payoff $\pi_{i}$ and $\pi_{j}$. With probability $1/\big{(}1\!+\!\exp[-30(\pi_{i}\!-\!\pi_{j})]\big{)}$, individual $i$ then reproduces and $j$ dies. Otherwise, individual $j$ reproduces and $i$ dies. During reproduction, there is a small chance that an offspring inherits an imperfect copy of the parent trait (with a mutation rate of $\mu\!=\!0.005$). In that case, the mutation is uniformly distributed around the parent trait in an interval of size 0.1. We iterate this simulation step a large number of times. At regular time intervals, we record the composition of the population, and the average cooperation rates and payoffs. For the simulations in Fig. 3, we consider comparably large populations ($10^{4}$ in the prisoner’s dilemma and the snowdrift game; $2\!\cdot\!10^{4}$ in the stag-hunt game).
Theoretical analysis
For our analytical results, we use the adaptive dynamics formalism [38, 39, 40]. Here, we assume infinitely large populations and rare mutations of infinitesimal magnitude. Under these assumptions we can focus on mutant traits close-by to the resident, such that $y\!\approx\!x$. A key quantity to study such mutants is the local fitness gradient or selection gradient $D(x)$, defined by
$D(x):=\left[\frac{\partial s_{x}(y)}{\partial y}\right]_{y=x}.$ (6)
The selection gradient determines the direction of evolutionary change. When $D(x)$ is positive (negative), mutants with a slightly higher (lower) trait value than $x$ replace the resident population. In that case, the resident trait value $x$ grows (decreases). A point $x^{*}$ that satisfies $D(x^{*})=0$ is called a singular point. A singular point is called convergence stable if
$\left[\frac{dD(x)}{dx}\right]_{x=x^{*}}\!<\!0.$ (7)
This condition determines whether the singular point can be reached at all, starting from a resident population sufficiently close-by. Moreover, the singular point is evolutionary stable if
$\left[\frac{\partial^{2}s_{x}(y)}{\partial y^{2}}\right]_{x=y=x^{*}}\!<\!0.$ (8)
In that case, the fitness function has a local maximum at $x^{*}$, and the singular point corresponds to an evolutionary endpoint. If the singular point is at a fitness minimum, however, this may indicate the presence of an evolutionary branching point. For Fig. 4, we verify the respective conditions numerically.
Data, Materials, and Software Availability. Simulations have been run with MATLAB. For our analytical results we used Mathematica. The respective simulation data and source code can be found in this OSF repository.
Author contributions. M.C.C., F.P.S. and C.H. designed research; M.C.C. performed research and analyzed data; and M.C.C., F.P.S. and C.H. wrote the paper.
Author declaration. The authors declare no competing interest.
Acknowledgments. M.C.C and C.H. acknowledge generous support by the European Research Council Starting Grant 850529: E-DIRECT. F.P.S acknowledges funding from the Dutch Research Council (NWO) through the project with file number OCENW.M.22.322 of the research programme Open Competitie ENW.